OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (8, -4) ; Slope: - 5/6

OpenStudy (anonymous):

A. y+4=-5/6(x+8) B. y-4=-5/6(x+8) C. y+4=-5/6(x-8) D. y-4=5/6(x-8)

OpenStudy (anonymous):

I am not good at slops but I thank c

OpenStudy (anonymous):

Four more

OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (1, 4) and has a slope of −3? A. y-1=-3(x-4) B. y+4=3(x+1) C. y+4=-3(x-1) D. y-4=-3(x-1)

OpenStudy (anonymous):

a